### 02 Mar 2024 v1

# MuseGraph: Graph-oriented Instruction Tuning of Large Language Models for Generic Graph Mining

Abstract.
Graphs with abundant attributes are essential in modeling interconnected entities and improving predictions in various real-world applications. Traditional Graph Neural Networks (GNNs), which are commonly used for modeling attributed graphs, need to be re-trained every time when applied to different graph tasks and datasets. Although the emergence of Large Language Models (LLMs) has introduced a new paradigm in natural language processing, the generative potential of LLMs in graph mining remains largely under-explored. To this end, we propose a novel framework MuseGraph, which seamlessly integrates the strengths of GNNs and LLMs and facilitates a more effective and generic approach for graph mining across different tasks and datasets. Specifically, we first introduce a compact graph description via the proposed adaptive input generation to encapsulate key information from the graph under the constraints of language token limitations. Then, we propose a diverse instruction generation mechanism, which distills the reasoning capabilities from LLMs (e.g., GPT-4) to create task-specific Chain-of-Thought-based instruction packages for different graph tasks. Finally, we propose a graph-aware instruction tuning with a dynamic instruction package allocation strategy across tasks and datasets, ensuring the effectiveness and generalization of the training process. Our experimental results demonstrate significant improvements in different graph tasks, showcasing the potential of our MuseGraph in enhancing the accuracy of graph-oriented downstream tasks while keeping the generation powers of LLMs.
Generic graph mining, Large language models, Instruction tuning
1. Introduction
Figure 1. An illustrative example of the need for a generic graph model that can be directly applied to various graph tasks and datasets.
Graphs with abundant attributes are widely used to model interconnected entities, and they are pivotal for enhancing downstream predictions across various real-world applications. The complexity and diversity of tasks across different datasets necessitate advanced models capable of harnessing graph information effectively.
Recently, Graph Neural Networks (GNNs) have become widely used in modeling attributed graphs (hamilton2017inductive, ; wang2019heterogeneous, ; cui2020adaptive, ). However, they need to be re-trained whenever applied to different graph datasets and tasks. Inspired by the great success of Large Language Models (LLMs), the combination of GNNs and LLMs aims to enhance the processing of text-attributed graphs, which can be categorized into two main approaches. The first line tries to train GNNs with LLM-enhanced features (e.g., OFA (liu2023one, ), ALL-in-One (sun2023all, ), GIMLET (zhao2023gimlet, ), ExpAsFeat (he2023explanations, ), LLMForGraph (chen2023exploring, ) and PRODIGY (huang2023prodigy, )) The second line tunes LLMs for graph applications (e.g., InstructGLM (ye2023natural, ), NLGraph (wang2024language, ) and LMMoL (qian2023can, ), GPT4GRAPH (guo2023gpt4graph, )).
Despite the promising direction of integrating LLMs with GNNs, the full exploration of LLMs’ generative capabilities has yet to be harnessed for a wide array of graph-oriented tasks and datasets. Figure 1 illustrates the necessity for a generic graph framework, which shows the multitude of task and dataset combinations possible. Such a generic graph model can not only capture the semantic and structural information of graphs for tasks like node classification and link prediction but also retain the rich generative power inherent in LLMs for tasks like graph-to-text applications. Such a model can also be adept across a spectrum of datasets (e.g., from MIMIC’s clinical data to arXiv’s academic texts), without sacrificing the generative abilities that are central to language models. However, three obstacles stand in the way of achieving this goal.
Challenge I: How to extract informative graph descriptions under the limitation of input token? To harness the full potential of Large Language Models for generic graph mining, an essential hurdle is translating the graph with abundant semantics and complex structures into a format that LLMs can process effectively, especially under strict input token limitations. Without a compact graph description that extracts key information from graphs within the LLMs’ token limitation, the model’s ability to grasp and utilize the graph’s semantic and structural richness is severely limited, potentially leading to suboptimal performance in graph applications.
Challenge II: How to automatically generate diverse instructions? Creating a diverse set of high-quality instructions for fine-tuning LLMs is essential for generic graph mining tasks. However, these instructions are often neither readily available nor economically feasible to produce manually. While advanced language models like GPT-4 possess the capability to generate diverse instructions, it remains unknown how to effectively leverage these generative abilities to produce relevant, task-specific instructions.
Challenge III: How to properly allocate instructions for graph-oriented instruction tuning? The effectiveness of instruction tuning for LLMs is significantly influenced by the selection of instructions, which enables the LLMs to comprehend and execute generic graph mining tasks effectively. However, maintaining broad task and dataset coverage while averting catastrophic forgetting poses a significant challenge, especially when dealing with a variety of tasks and datasets in one generic graph model.
To tackle these challenges, we propose Graph-oriented Instruction Tuning of Large Language Models for Generic Graph Mining (MuseGraph), which consists of three pivotal steps: (i) Development of Compact Graph Descriptions, where we introduce a novel “node energy” metric to texualize graphs with essential semantic and structural details under limited tokens; (ii) Generation of Diverse Instructions, which invokes the advanced generative abilities of GPT-4 to create CoT-based instruction packages tailored for various graph tasks, thus enriching LLMs’ capabilities in understanding and analyzing graph data without the expense of manual instruction crafting; (iii) Graph-aware Instruction Tuning, which introduces a dynamic instruction package allocation strategy based on the specific needs of each graph mining task, ensuring comprehensive and effective LLM tuning.
Our overall contributions in this work are summarized as follows:
  • Formulation of generic graph mining. We establish a generic graph framework that effectively transforms graph structures into LLM-friendly formats while preserving the generative capabilities necessary for diverse graph-oriented tasks
  • Effective Model Designs. We design and implement a set of models and mechanisms including the development of compact graph descriptions, automatically generating diverse, task-specific instructions, and graph-aware instruction tuning, targeting a generic graph model across tasks and datasets.
  • Extensive Experiments Across Graph Tasks and Datasets. We conduct thorough experiments to validate our approach, demonstrating its superiority over existing methods in a variety of graph tasks and datasets, highlighting its effectiveness and interpretability in enhancing LLMs for graph mining.
2. Related Work
2.1. Semantic-rich Graph Representation Learning
Graph Representation Learning has emerged as a key technique for the complex structures of networks with abundant attributes (DBLP:journals/debu/HamiltonYL17, ; DBLP:journals/tkde/CaiZC18, ; DBLP:journals/tnn/WuPCLZY21, ; zhou2020graph, ; DBLP:journals/tbd/WangBSFYY23, ). Among existing node embedding methods, many had analyzed and utilized the great promise of random walks (RWs) in capturing the topological structures of graphs (perozzi2014deepwalk, ; grover2016node2vec, ; dong2017metapath2vec, ; ivanov2018anonymous, ). However, the above methods ignored the abundant attribute information surrounding the nodes and edges (mao2023hinormer, ). Recently, Graph neural networks (GNNs) for learning node representations through aggregating information from neighboring nodes on graphs (kipfsemi, ; hamilton2017inductive, ; yang2020relation, ). However, most existing GNNs were established in a supervised learning setting, which requires abundant task-specific labeled data that may not be available in real-world applications (zhu2021graph, ; cooray2022graph, ), and the embeddings they learned are not generalizable across different downstream tasks (zhang2018link, ), which had to be re-trained whenever applied to different graph tasks and dataset.
Although some studies tried to reduce the labeling effort by pre-training an expressive GNN model on unlabeled data with self-supervision methods (e.g., contrastive learning) (hu2020gpt, ; zhu2021transfer, ; jiang2021pre, ), their performances in specific downstream tasks still relied much on the properly chosen self-supervision tasks and attribute encoders (zhang2023structure, ; sun2022gppt, ). Therefore, there is still a lack of a uniform framework for generic graph mining across different tasks and datasets.
2.2. Leveraging LLMs for Graph Mining
Figure 2. The overall framework of MuseGraph.
Table 1. A comparison between MuseGraph and related methods, where GNN* denotes the method with GNN training and LLM* denotes the one with LLM fine-tuning. The partial checkmark $\sim$  for some methods means that they can be extended to achieve the targets, but this is not their main research focus.
GNN* LLM* Cross-task Cross-dataset
ExpAsFeat (he2023explanations, )
PRODIGY (huang2023prodigy, )
LLMForGraph (chen2023exploring, )
All-in-One (sun2023all, )
OFA (liu2023one, )
NLGraph (wang2024language, ) $\sim$
GPT4GRAPH (guo2023gpt4graph, ) $\sim$
GraphGPT (guo2023gpt4graph, ) $\sim$
InstructGLM (ye2023natural, ) $\sim$
MuseGraph
Motivated by the impressive strides in Large Language Models (LLMs), combining Graph Neural Networks (GNNs) and LLMs is creating substantial progress in handling complex text-attributed graphs (li2024survey, ; chen2023label, ; chen2023exploring, ). Methods leveraging this combination fall into two broad categories (shown in Table 1).
The first category is training GNNs with LLM-enhanced features, labeled GNN* in Table 1). For example, ExpAsFeat (sun2023all, ) and PRODIGY (huang2023prodigy, ) utilized LLMs to encode the textual information associated with nodes on the graph. LLMForGraph (chen2023exploring, ) enabled LLMs to improve text-attributed graph processing by deepening semantic understanding and leveraging extensive knowledge bases. All-in-One (sun2023all, ) combined NLP prompting concepts into graph downstream tasks. OFA (liu2023one, ) prompted LLMs to translate diverse textual attributes into uniform feature vectors across domains.
These models primarily utilized LLMs to process textual content within graphs and tend to be tailored to specific domains, often failed to be applied for cross-task and cross-dataset applications and lost the generation capability from LLMs. Although these models effectively focus on leveraging LLMs’ textual prowess, most of them are domain-specific graph models, which fail to perform cross-task and cross-dataset.
The second category denoted as LLM* in Table 1, mainly focuses on tuning LLMs for graph-related tasks. For example, InstructGLM (ye2023natural, ), NLGraph (wang2024language, ), and GPT4GRAPH (guo2023gpt4graph, ) leveraged LLMs to understand and analyze graph data through textual descriptions. GraphGPT (tang2023graphgpt, ) with one checkmark on GNN* and a partial checkmark on LLM* jointly trained task-specific GNN and one projector on the top of LLMs, which performed well across different datasets. Note that, the partial checkmark for some methods means that they can be extended to cross-dataset targets, but this is not their main research focus and they have no design mechanism to achieve it. Compared with GNN*-based methods, LLM*-based methods have the potential to perform generic graph mining with powerful cross-tasks and cross-datasets. However, such power remains largely under-explored. Table 1 summarizes the characteristics of different models. We can see that our MuseGraph is the most generic model among existing works.
3. The MuseGraph Framework
3.1. Overview of Our Framework
Objective: In this paper, we aim to pursue an ideal graph that can fascinate a more effective and generic approach via graph-oriented instruction tuning of LLMs, through which we wish to seamlessly integrate the strengths of GNNs and LLMs for graph mining across different tasks and dataset
Overview: To achieve this goal, we propose MuseGraph framework, which comprises three major components. Firstly, we develop a compact graph description mechanism that captures critical semantic and structural details within the constraints of language tokens. Secondly, we generate a diverse range of instructions through the reasoning capabilities of LLMs, thus facilitating task-specific, Chain-of-Thought-based instruction packages for various graph tasks. Thirdly, we adopt a graph-aware instruction tuning, utilizing a dynamic allocation strategy for instruction packages tailored to the unique demands of each graph task. The overall model architecture is shown in Figure 2, and we elaborate on the main components.
3.2. Compact Graph Description
Leveraging the capabilities of Large Language Models (LLMs) for graphs presents a unique set of challenges, primarily due to LLMs’ inherent limitations in directly inputting graph structures under the token limitation. Therefore, it is non-trivial to automatically extract key information from graphs under the token limitations, which require a compact description including complex node and edge attributes along with structural details.
Inspired by common graph analysis techniques such as walks and neighbors, we propose a novel method of textualization to describe graphs via these concepts. In this way, neighbors are helpful to understand local connectivity and feature distribution (kipfsemi, ; hamilton2017inductive, ; yang2020relation, ), providing a granular view of node attributes; while walks offer a dynamic method to explore the graph’s structure and the relationships between nodes, highlighting the diversity of connectivity and paths (micali2016reconstructing, ; ivanov2018anonymous, ). The integration of neighbors and walks can achieve a holistic understanding of graph structures.
Given LLM token limits and the varied contribution of neighbors and walks to node understanding, we further develop an adaptive input generation mechanism to ensure the compactness of the description. Specifically, we first design a “node energy” metric, $H(v_{i})$ assessing node information from two perspectives: token count in attributes and node degree count. This metric enables us to effectively filter and select neighbor nodes and walk nodes, prioritizing those that are abundant in semantic information and possess a significant number of neighboring nodes, thus enhancing the expressiveness of the graph description. The calculation of node energy $H(v_{i})$ is formulated as follows:
(1) $H(v_{i})=T(v_{i})*\lceil\log(D(v_{i})+1)\rceil,$
where $v_{i}$ is the target node, ${T(v_{i})}$ is the number of node tokens processed by a tokenizer and ${D(v_{i})}$ is the number of node degrees.
10:  Attributed graph $\mathcal{G}$ with $N$ node, token count set $\mathcal{T}$,
node energy set $\mathcal{H}$, target node $v^{*}$, token limitation $L(v^{*})$
0:  Key neighbor set $\mathcal{N}(v^{*})$, key walk set $\mathcal{W}(v^{*})$
1:  Initialize $\mathcal{N}(v^{*})$, $\mathcal{W}(v^{*})$ as empty
22:  Select $v_{i}\in G(v^{*})$ with $H(v_{i})\geq H(v^{*})$ and $L(v^{*})\geq T(v_{i})$
for $\mathcal{N}(v^{*})$, where $G(v^{*})$ is $v^{*}$’s one-hop neighbors
33:  Expand $\mathcal{W}(v^{*})$ starting from $v^{*}$ based on $\mathcal{G}$ within $L(v^{*})$
and $H(v^{*})$ constraints
Algorithm 1 Adaptive Input Generation
Figure 3. An example illustrates how to extract the key information of nodes from neighbors and walks based on node energy $H(v_{i})$ and the limitation of language tokens $L(v_{i})$.
Based on $H(v_{i})$, our method strategically incorporates neighbors and walks, as depicted in Algorithm 1. A neighbor, $v_{i}$, is chosen to describe the target $*v*$ if its $H(v_{i})$ surpasses $H(v^{*})$, ensuring the included node can provide supplemental information. The process of including walks concludes when encountering a node whose $H(v_{i})$ does not meet the threshold set by the target’s $H(v^{*})$, thus refining the input to maximize relevance within the constraints of the token limit. To further demonstrate the adaptive input generation for neighbors and walks tailored to each node, we present an example in Figure 3). Node $v_{1}$’s with a relatively low $H(v_{1})$ value, includes a wide range of neighbors to capture more context, excluding $v_{4}$ due to its even lower $H(v_{4})$. Given the token limitations, only two walks are sampled for $v_{1}$ to maintain a compact input. Conversely, $v_{2}$ with its higher $H(v_{2})$, inherently carries more information, prompting the selection of fewer neighbors. $v_{1}$ is excluded since $H(v_{1})<H(v_{2})$. This frees up tokens to detail more walks for $v_{2}$. The $H$ metric thus finely balances neighbor and walk inclusion for each node, marrying information-rich and token-efficient. The process culminates in the textualization of each node’s key information, producing a tailored and compact graph description as depicted in the upper right of Figure 2. Note that, a graph-related task can involve multiple nodes. In this case, we further allocate each node to satisfy the token limit requirement by calculating the ratio of all node energies using the softmax function.
3.3. Diverse Instruction Generation
With the compact graph descriptions of each node, another necessary step in fine-tuning LLMs for generic graph mining lies in crafting diverse and high-quality instructions (ouyang2022training, ; jiao2023instruct, ). However, such well-constructed instructions are not readily available in every situation and manual construction is often very costly.
To address these problems, we propose to distill generation capabilities from advanced LLMs (shridhar2023distilling, ) (e.g., GPT-4 (achiam2023gpt, ) with over 200 billion parameters) for graph-related tasks. Our approach, inspired by the Chain-of-Thought (CoT) processing methodology (wei2022chain, ), prompts GPT-4 via a flexible template based on our compact graph descriptions for different tasks, and then constructs task-specific CoT-based instruction packages. Different from the existing methods that leverage CoT in the prompting stage (wei2022chain, ; xiang2024badchain, ), we directly construct CoT-based instruction packages, which can leverage its diversity and reasoning ability to facilitate instruction tuning with graph data.
Figure 4. A process showing how to distill capabilities from GPT-4 via Chain-of-Thought with node classification.
Specifically, we first design diverse CoT templates for targeted popular graph tasks, such as node classification, link prediction, and graph-to-text (shown in A Table 7). Then, we prompt GPT-4 to generate a small number of CoT-based instructions across diverse tasks. This approach distills GPT-4’s vast knowledge base to augment the reasoning and analytical abilities of our MuseGraph, as depicted in Figure 4.
To optimize the cost-effectiveness of querying GPT-4, we introduce the CoT-based instruction package. For every set of 1,000 standard instructions, we integrate 100 CoT-based instructions tailored to the same graph task. This approach not only proves to be economical but also broadens the diversity and flexibility of instructions, accommodating an array of graph-related tasks.
3.4. Graph-aware Instruction Tuning
The effectiveness of instruction tuning for LLMs is significantly influenced by the selection of instructions, which enables the LLMs to comprehend and execute generic graph mining tasks effectively (wang2023far, ; wang2022self, ). However, maintaining broad task and dataset coverage while averting catastrophic forgetting (mccloskey1989catastrophic, ; mosbach2020stability, ; vijay2022nerda, ), where LLMs might regress in their generative abilities, poses a significant challenge.
To this end, we propose a dynamic instruction package allocation strategy to adjust the volume of task-specific CoT-based instruction packages based on the complexities of tasks and datasets, which ensures that more complex tasks/datasets receive a proportionally larger set of instructions for detailed guidance. The calculation process unfolds in two phases:
For task complexity, we count the average output tokens for each task, reflecting the task-specific demands within a dataset. For dataset complexity, we compute the total node energy $H(v_{i})$ (cf. Eq. 1) of each graph data, using this metric to further fine-tune the instruction distribution. By performing a precise context-aware allocation of instructions, we can enhance the LLMs’ ability to learn and retain abundant knowledge across a diverse range of graph mining challenges.
To achieve the balance between effectiveness and efficiency, we adopt a graph-aware instruction tuning mechanism, which can sufficiently utilize diverse and high-quality instructions for further fine-tuning LLMs (ouyang2022training, ). Specifically, we adopt a general LLM LLaMA-V1-7B (abbr. LLaMA) (touvron2023llama, ) with LoRA (hu2021lora, ) as our starting point for fine-tuning. Then, based on diverse instruction set =$\mathcal{I}=\{I_{1},I_{2},\dots,I_{N}\}$ as model input, we adopt the negative log-likelihood loss as the fine-tuning objective (achiam2023gpt, ; ye2023natural, ), which is formulated as follows:
(2) $p_{\theta}\left(Y_{j,k}|I_{j},Y_{j,<k}\right)=LLM_{\theta}\left(I_{j},Y_{j,<k}\right),$
(3) $\mathcal{L}_{\theta}=-\sum_{k=1}^{|Y_{j}|}{\log p_{\theta}\left(Y_{j,k}|I_{j},Y_{j,<k}\right)},$
where $\theta$ is the learnable parameters of our graph-aware LLM, $I_{j}\in\mathcal{I}$ is the input sentence of LLM, and $Y_{j}$ is the output sentence of LLM.
After obtaining the fine-tuned LLM for generic graph mining (i.e., MuseGraph), we can apply it to various downstream tasks, such as node classification, link prediction, and graph-to-text.
Model Extension. Our framework is a fundamental approach to integrate LLMs and graph data for generic graph models, which can be based on different LLMs (e.g., Baichuan-7B (yang2023baichuan, ) and vicuna-7B-v1.1 (chiang2023vicuna, )). Moreover, we can adopt different parameter-efficient training approaches (e.g., QLoRA (dettmers2023qlora, ) and AdaLoRA (zhang2023adaptive, )). A full exploration of different LLMs is orthogonal to the main contributions in this work, which is left as future work.
Datasets Extension. Our method introduces a generic graph framework to simultaneously capture the semantic and structural information of attributed graphs across tasks and datasets. This allows us to leverage the capabilities of LLMs for genetic graph learning. Note that, our method only requires some meaningful attributes on the graphs, which are available in most real-world graphs such as biological networks, social networks, and knowledge graphs.
4. Experiment
In this section, our research undertakes multiple experiments to confirm MuseGraph framework’s efficiency in diverse conditions, addressing essential research inquiries:
  • RQ1: How does MuseGraph perform in comparison to state-of-the-art graph-oriented methods for generic graph mining?
  • RQ2: What are the effects of different model components?
  • RQ3: Can MuseGraph provide the generation capability and interpretability for generic graph mining?
4.1. Experimental Setup
4.1.1. Datasets Descriptions.
To comprehensively verify the effectiveness of our method, we use three real-world datasets for node classification, i.e., OGB-arxiv (abbr. Arxiv) (hu2020open, ), MIMIC-III (johnson2016mimic, ) and Cora (mccallum2000automating, ), and two for graph-to-text, i.e., AGENDA (koncel2019text, ) and WebNLG (gardent2017creating, ). The detailed statistics are shown in Table 2 and Table 3. To verify the generalizability and adaptability of models in node classification, we perform dataset splits for node classification following different ratios of 5:1:4 for Arxiv, 7:1:2 for MIMIC-III, and 2:1:7 for Cora. AGENDA and WebNLG are partitioned into training, validation, and testing sets with a 7:1:2 ratio. For more detailed descriptions of datasets, please refer to Appendix A.1.
4.1.2. Evaluation Protocols.
We evaluate the node classification performance using three commonly adopted evaluation metrics: Macro-F1 (tan2023walklm, ), Micro-F1 (tan2023walklm, ), and Weighted-F1 (fu2023hyperbolic, ); while evaluating graph-to-text performance using four evaluation metrics: BLEU-4 (papineni2002bleu, ), METEOR (banerjee2005meteor, ), ROUGE-L (lin2004rouge, ), and CHRF++ (popovic2017chrf++, ). The F1 score is a metric of the model’s accuracy in binary and multi-class classification tasks, which considers both precision and recall. BLEU-4 and ROUGE-L compute the ratios of overlapping and matching between generated and real text. METEOR computes the harmonic mean of precision and recall. CHRF++ computes F-score averaged on both character and word-level n-grams.
Table 2. Statistics of the datasets for node classification.
Dataset
Arxiv MIMIC-III Cora
# Nodes
169,343 32,267 25,120
# Edges
1,116,243 559,290 182,280
# Label nodes
169,343 4,880 17,093
# Classes
40 19 70
Table 3. Statistics of the datasets for graph-to-text.
Dataset
AGENDA WebNLG
# Graphs
40,720 19,945
# Relations
7 373
Avg. # Nodes
12.37 5
Avg. # Triples
4.48 4
Avg. Length
140.36 24.21
Table 4. Experimental results on three benchmark datasets for node classification. The best performances are highlighted in boldface and the second runners are underlined. MuseGraph achieves the best performance on both datasets.
Method Macro-F1 Micro-F1 Weighted-F1 Macro-F1 Micro-F1 Weighted-F1 Macro-F1 Micro-F1 Weighted-F1
Dataset Arxiv MIMIC-III Cora
MLP 0.2376 0.4048 0.3987 0.3079 0.3996 0.3480 0.0854 0.1213 0.1173
GraphSAGE 0.3025 0.5252 0.5169 0.5227 0.5861 0.5614 0.0927 0.1518 0.1418
GCN 0.3158 0.5119 0.5009 0.5396 0.5768 0.5446 0.0902 0.1479 0.1459
GAT 0.2997 0.4983 0.4878 0.6193 0.6452  0.6331 0.0879 0.1480 0.1432
RevGNN 0.3049 0.5182 0.5071 0.5419 0.5733 0.5571 0.0913 0.1488 0.1469
DGI 0.2792 0.4443 0.4348 0.4657 0.4931 0.4748 0.0785 0.1278 0.1256
GKD 0.2843 0.4513 0.4402 0.4593 0.4826 0.4671 0.0450 0.1577 0.1358
GLNN 0.3542 0.6025 0.5889 0.5346 0.5765 0.5527 0.1059 0.1729 0.1628
NodeFormer 0.3082 0.5321 0.5203 0.5612 0.5861 0.5795 0.0938 0.1542 0.1539
DIFFormer 0.3195 0.5374 0.5252 0.5792 0.5954 0.5907 0.0946 0.1518 0.1436
BART 0.3285 0.5267 0.5191 0.5787 0.5896 0.5822 0.0989 0.1581 0.1543
T5 0.3147 0.5309 0.5226 0.5681 0.5983 0.5873 0.0955 0.1567 0.1522
LLaMA 0.3159 0.5274 0.5182 0.5624 0.5989 0.5881 0.0986 0.1584 0.1545
MuseGraph 0.4245 0.6920 0.6773 0.6875 0.7275 0.7168 0.1197 0.2013 0.2109
Table 5. Experimental results on two benchmark datasets for graph-to-text. The best performances are highlighted in boldface and the second runners are underlined.
Method BLEU-4 METEOR ROUGE-L CHRF++ BLEU-4 METEOR ROUGE-L CHRF++
Dataset AGENDA WebNLG
GraphWriter 0.1413 0.1892 0.2761 0.3835 0.4584 0.4021 0.6062 0.5543
CGE-LW 0.1801 0.2234 0.2562 0.4679 0.4860 0.4345 0.6252 0.5856
BART 0.2365 0.2519 0.2876 0.5044 0.5249 0.4223 0.6561 0.7204
T5 0.2215 0.2373 0.3068 0.4872 0.5878 0.4418 0.6824 0.7439
GPT-3.5 0.1057 0.1702 0.2522 0.4586 0.1108 0.2389 0.3587 0.4875
MuseGraph 0.2231 0.2457 0.3098 0.4913 0.5359 0.4386 0.6973 0.7276
4.1.3. Implementation Details.
All the compared baselines are optimized through the Adam optimizer and the learning rate is searched in [1e-4, 1e-2]. The hyper-parameters of baselines are chosen carefully based on either grid search or their official source codes. The training process of instruction-tuning is carried out for one epoch.
4.1.4. Methods for Comparison
In our experiment, we consider various state-of-the-art methods for comprehensive evaluation. For node classification, we compare our proposed MuseGraph with the following 13 representative baselines from two main perspectives:
For graph-to-text, we compare MuseGraph with the following 5 baselines from two perspectives:
For more details of the compared baselines, refer to Appendix A.2.
4.2. Overall Performance Comparison (RQ1)
To evaluate our MuseGraph framework, we conducted extensive testing across node classification and graph-to-text tasks using five real-world datasets.
4.2.1. Main Results Compared with GNN-based Methods.
MuseGraph stands out with its single model finely tuned across the Arxiv, MIMIC-III, Cora, AGENDA, and WebNLG datasets. This answers RQ1, showcasing remarkable cross-task and cross-dataset generalization capabilities. Moreover, our generic MuseGraph is in contrast to most GNN-based methods that typically require dataset-specific training, such as GAT and GLNN for node classification while GraphWriter and CGE-LW for graph-to-text.
In the node classification task, MuseGraph demonstrates significant performance gains, ranging from reasonably large (12.33% over GAT on MIMIC-III) to significantly large (19.67% over GLNN on Cora). In the graph-to-text task, the performance gains of MuseGraph over all GNN-based baselines range from 1.02% (achieved in METEOR on WebNLG) to 57.89% (achieved in BLEU-4 on AGENDA). Note that, GNN-based methods (e.g., GLNN and GAT) only achieve second-runner in the node classification task. Since the generative and reasoning abilities of GNN-based methods are weaker than LLM-based ones in generative tasks, they can not achieve the second-runner in the graph-to-text task. In particular, GLNN is optimized for attribute-rich datasets (e.g., Arxiv and Cora) through its knowledge distillation mechanisms, while GAT is tailored for datasets with basic attributes (e.g., MIMIC-III) through its ability to weigh nodes differently. Compared with GLNN and GAT, our MuseGraph with graph-ware instruction tuning with the diverse CoT-based instruction packages can effectively comprehend graph data and achieve accurate predictions of various downstream tasks.
One step further, since one key challenge of node classification lies in the generalizability and adaptability of models, we design a few-shot setting on Cora to evaluate models in extending knowledge to unseen scenarios and adapting to new tasks with limited training data. As shown in Table  4, our MuseGraph can stay strong with a small size of training data, where we win the performance gains over the second-best performance ranging from 13.03% in Macro-F1 to 29.55% in Weighted-F1. Through the dynamic instruction package allocation strategy across tasks and datasets, our proposed MuseGraph can understand and analyze the meaningful graph attributes and structures, which can help MuseGraph adapt to different graphs. Consequently, MuseGraph maintains superior performance even in the few-shot setting.
4.2.2. Main Results Compared with LLM-based Methods.
To fairly compare our MuseGraph with the other LLM-based methods, we fine-tune LLMs like BART, T5, and LLaMA with LoRA to better adapt them to our target graph mining tasks. In node classification tasks, these LLM-based methods demonstrate stable performance across various tasks and datasets. However, their predictive capabilities are primarily confined to text-based information. With the designed compact graph description, diverse instruction generation, and graph-aware fine-tuning mechanism, MuseGraph surpasses these LLM-based methods by harnessing both textual and structural graph information, leading to a more comprehensive understanding and enhanced performance (with up to 38.57% compared with the second-runner).
Compared with GNN-based methods, BART and T5 exhibit superior performance in graph-to-text, reaching state-of-the-art on certain metrics (e.g., BART on AGENDA and T5 on WebNLG). The success of BART and T5 is attributed to their sophisticated pre-training on diverse text corpora, which equips them with a comprehensive understanding of complex knowledge graph structures and precise word usage. MuseGraph, with less pre-training data towards graph-to-text task and mainly targeting generic graph mining, still retains competitive performance against these models by effectively leveraging our graph-oriented instruction tuning of LLMs.
Note that, MuseGraph can sometimes exceed all baselines with ROUGE-L metric, achieving an average performance gain of 22.40% over the second-runner T5 on average. By extracting the compact graph description by adaptively allocation the neighbors and walks under the input token limitation, MuseGraph adeptly extracts a compact graph description. This enables it to accurately grasp both the semantic and structural nuances of graphs through a graph-aware fine-tuning mechanism. Given that ROUGE-L prioritizes the longest common subsequence between generated and reference texts, MuseGraph’s focus on key graph information translates into texts that are both accurate and coherent. This capability is particularly valued in ROUGE-L assessments, which do not require strict adherence to the sequence of words and phrases at the word or phrase level. As a result, MuseGraph’s generated texts can achieve high scores even when the structure is more flexible.
Figure 5. Comparison of results for node classification on Arxiv within different allocations of CoT-based instruction packages across various tasks and datasets.
Table 6. Example graph-to-text results generated from LLaMA, GPT-3.5 and MuseGraph on AGENDA. The output includes the generated abstract and the corresponding Chain-of-Thought. Best viewed in color.
Input: The compact graph description of this PAPER is listed as follows: Ego graph nodes: {1. METHOD: [Incremental fuzzy Learning Algorithm, Fuzzy Intelligent System]; 2. OTHERSCIENTIFICTERM: [Priority Values, Fuzzy Rule]; 3. TASK: [Automotive Engineering Diagnosis]} Random walks: {1. Incremental fuzzy Learning Algorithm USED-FOR Automotive Engineering Diagnosis. 2. Fuzzy Intelligent System USED-FOR Automotive Engineering Diagnosis. 3. Priority Values USED-FOR Incremental fuzzy Learning Algorithm USED-FOR Fuzzy Intelligent System USED-FOR Automotive Engineering Diagnosis. 4. Fuzzy Rule USED-FOR Fuzzy Intelligent System USED-FOR Automotive Engineering Diagnosis.}
Instruction: Given the target PAPER Incremental Learning in a Fuzzy Intelligent System with the compact graph description. Please generate the target PAPER’s abstract and the Chain-of-Thought process that could lead to this generation.
LLaMA: In this paper, we propose an incremental fuzzy learning algorithm for a fuzzy intelligent system. The incremental fuzzy learning algorithm is based on the priority values of fuzzy rules and the fuzzy belief values of the fuzzy rules. The incremental fuzzy learning algorithm is applied to an automotive engineering diagnosis. The CoT for this generation is as follows: 1. Focus: Fuzzy learning algorithm for diagnosis. 2. Function: Uses priority values for decision-making. 3. Incrementality: Adjusts based on new data, and maintains old knowledge. 4. Usage: Part of a fuzzy system for vehicle diagnostics. 5. Outcome: Aimed at better diagnosis through incremental learning and data emphasis.
GPT-3.5: This paper presents a novel approach to incremental learning within a fuzzy intelligent system framework tailored for automotive engineering diagnosis. We introduce an Incremental Fuzzy Learning Algorithm designed to enhance the adaptability and efficiency of diagnostic systems in automotive engineering contexts. Leveraging fuzzy logic principles, our approach allows for continuous learning and refinement of diagnostic capabilities over time. Central to our methodology is the integration of priority values and fuzzy rules, which facilitate the dynamic adjustment of learning priorities and the refinement of fuzzy inference mechanisms. Through extensive experimentation and evaluation, we demonstrate the effectiveness and practical utility of our approach in enhancing the accuracy and reliability of automotive engineering diagnosis systems. Our findings underscore the importance of incremental learning techniques within fuzzy intelligent systems for advancing diagnostic capabilities in complex engineering domains. The CoT for this generation is as follows: 1. Main method: Incremental fuzzy learning algorithm. 2. Purpose: Improve automotive engineering diagnosis. 3. Components: Prioritizes fuzzy belief values, priority values, and rules. 4. Learning process: Algorithm learns incrementally, adapting to new data. 5. System integration: Used within a fuzzy intelligent system for diagnostics.
MuseGraph: This paper describes an incremental fuzzy learning algorithm for a fuzzy intelligent system that is used in automotive engineering diagnosis. The incremental fuzzy learning algorithm uses priority values to determine the fuzzy rules that should be updated. The incremental fuzzy learning algorithm also uses fuzzy belief values to determine which fuzzy rules should be updated. The incremental fuzzy learning algorithm is incremental in that incremental fuzzy learning algorithm only updates the fuzzy rules that have changed. The incremental fuzzy learning algorithm is also incremental in that the incremental fuzzy learning algorithm only updates the fuzzy rules that are necessary to maintain the fuzzy intelligent system’s performance. The CoT for this generation is as follows: 1. Identify main concept: Incremental fuzzy learning algorithm within a fuzzy intelligent system 2. Clarify task relevance: Algorithm is used for automotive engineering diagnosis. 3. Detail key attributes: Algorithm utilizes fuzzy belief values and priority values. 4. Define algorithm benefits: Detects and updates from new data, retains useful knowledge, emphasizes new learning. 5. Apply to system: The fuzzy intelligent system incorporates the algorithm for enhanced diagnostic capabilities.
4.3. Ablation Studies (RQ2)
To study the effectiveness of different allocations of CoT-based instruction packages across various tasks and datasets, we conduct ablation studies for node classification on Arxiv, which is based on our proposed compact graph description for inputting graph. Using LLaMA as our baseline model, the studies evaluated our MuseGraph across various configurations, while maintaining an equal total count of instruction packages. We study our MuseGraph as follows:
  • 1 Task+1 Dateset: Utilizes our proposed CoT-based instructions for node classification on Arxiv.
  • 2 Task+1 Dateset: Incorporates instructions from both node classification and link prediction on Arxiv.
  • 2 Task+2 Dateset: Expands to include node classification and link prediction instructions from both Arxiv and Cora.
  • 3 Task+3 Dataset: Engages with instructions across node classification, link prediction on Arxiv and Cora, and graph-to-text on AGENDA..
As shown in Figure 5, we have the following observations:
Compared with 1 Task+1 Dataset and LLaMA, 2 Task+1 Dataset leads to performance gains ranging from 3.22% (achieved in Micro-F1) to 18.30% (achieved in Macro-F1). Such results show the effectiveness of diverse task-specific CoT-based instruction packages within the token limitation for instruction tuning, which can enable the LLMs to comprehend and execute generic graph mining tasks.
Furthermore, with increasing the number of datasets, the performance gains of 2 Task+2 Dataset over 2 Task+1 Dataset ranges from 4.96% (achieved in Micro-F1) to 7.24% (achieved in Weighted-F1). This indicates that reasonable allocation of the different volumes of CoT-based instruction packages across datasets can help LLMs effectively and efficiently perform predictions of graph-related tasks.
With adding more types of tasks and datasets, our proposed MuseGraph can achieve further improvements, where 3 Task+3 Dataset overperforms 2 Task+2 Dataset ranges from 3.03% in Macro-F1 to 3.94% in Weighted-F1. Such observation strongly indicates that the dynamic instruction package allocation strategy can enhance the accuracy of graph tasks and avert catastrophic forgetting, where LLMs might regress in their generative abilities.
4.4. Case Studies (RQ3)
To demonstrate the advantages of MuseGraph over the LLM-based methods on the generation tasks, we provide the abstract of graph-to-text task prompted by LLaMa, GPT-3.5, and our MuseGraph on AGENDA. The provided graph structure and its ground-truth abstract are displayed at the top of Table 6. We standardized the process by employing our compact graph description as the input across all models.
As shown in Table 6, the abstract generated by LLaMA offers a basic summary while overlooking some key details, such as the specific incremental learning aspect of the fuzzy algorithm (shown in the blue text). Compared with LLaMA, Chatgpt-3.5 generates even three times longer content with an elaborate description. However, the model inevitably includes some fabricated sentences, such as the “Through extensive experiment and evaluation, …” in red, which is irrelevant to the provided ground-truth abstract.
Since our method performs graph-aware instruction tuning with compact graph descriptions, we can produce precise abstracts by extracting the key information from the graph. Moreover, by incorporating the distilled reasoning and generative capacity from CoT-based instructions, MuseGraph can not only produce an abstract that aligns with the core details of the ground-truth but also provides fluent and precise CoT inherited the advanced large language models.
5. Conclusion
In this paper, we introduce MuseGraph, an effective and generic approach for graph mining across different tasks and datasets, which can also alleviate the challenges posed by language token limitations. Through the innovative design of compact graph descriptions with adaptive input generation, the generation of diverse, task-specific instructions, and the implementation of graph-aware instruction tuning, our MuseGraph integrates the strengths of Graph Neural Networks (GNNs) and Large Language Models (LLMs). Our comprehensive experimental results demonstrate MuseGraph’s superior performance against state-of-the-art baselines across graph tasks and datasets, illustrating its ability not only to enhance the precision of graph-oriented downstream tasks but also to preserve the generative capabilities of LLMs, which is further consolidated with our real case study results.
Appendix A Appendix
A.1. Detailed Descriptions of Datasets
This section provides detailed descriptions of each graph dataset used in our experiment.
  • OGB-arxiv (abbr. Arxiv) represents a directed graph that captures the citation network among computer science arXiv papers indexed by MAG [58]. Each paper in the dataset is associated with a research category, manually labeled by the authors and arXiv moderators. These research categories are selected from a set of 40 subject areas.
  • MIMIC-III is a graph of diseases, patients, and visits, where nodes and relations are extracted from clinical records [26]. Diseases are classified into 19 categories according to ICD-9-CM.
  • Cora is known as the “Cora Research Paper Classification” dataset. It is a collection of research papers that are linked to each other through citations. The abstract of a paper is deemed a text document. The papers are classified into a topic hierarchy with 73 leaves. We utilize an expanded version, which is larger and has more classes (70 in total) compared to previous versions [27].
  • AGENDA (Abstract Generation Dataset) is a dataset that pairs knowledge graphs with paper abstracts from scientific domains. The graphs in AGENDA were automatically extracted from the SciIE information extraction system [35]. Each instance in AGENDA includes the title, entities, graph, and abstract of a paper. We use the entities and graphs as input for the models.
  • WebNLG is a crowd-sourced RDF triple-to-text dataset manually crafted by human annotators. The dataset contains graphs from DBpedia [2] with up to 7 triples paired with one or more reference texts. We take data from WebNLG v2.1 for experiments.
A.2. Details of Compared Baselines
  • MLP [71] The method employs a Multi-layer Perceptron for node representation.
  • GraphSAGE [16] is a general, inductive framework that leverages node feature information, such as text attributes to efficiently generate node embeddings.
  • GCN [27] scales linearly in the number of graph edges and learns hidden layer representations that encode both local graph structure and features of nodes.
  • GAT [54] novel neural network architectures that operate on graph-structured data, leveraging masked self-attentional layers to address the shortcomings of prior methods based on graph convolutions or their approximations.
  • RevGNN [31] is designed to enable the capture of long-range interactions in graph-structured data without suffering from the vanishing gradient issue, due to its reversible computation framework that preserves information flow throughout the layers.
  • DGI [55] is a general approach for learning node representations within graph-structured data in an unsupervised manner.
  • GKD [69] is a new paradigm of knowledge transfer that aims at encoding graph topological information into graph neural networks (GNNs) by distilling knowledge from a teacher GNN model trained on a complete graph to a student GNN model operating on a smaller or sparser graph.
  • GLNN [75] is a combination of GNNs and MLPs through knowledge distillation (KD).
  • NodeFormer [65] is a scalable graph structure learning transformer for node classification.
  • DIFFormer [64] is a scalable graph transformer induced by energy-constrained diffusion.
  • GraphWriter [29] uses Graph Neural Networks (GNNs) and transformer-based models to produce coherent texts from structured graph inputs.
  • CGE-LW [47] proposes a graph-to-text model by combining both global and local node aggregation strategies.
  • BART-large [30] linearizes the KG into sequence and applies BART-large to generate text.
  • T5-large [46] linearizes KG into a triple sequence and employs T5-large to generate text.
  • LLaMA-V1-7B [53] is an open-sourced Large Language Model (LLM) designed for natural language understanding and generation tasks.
  • ChatGPT-3.5 [41] uses Graph Neural Networks (GNNs) and transformer-based models to generate coherent text based on input structured graphs.
A.3. Instruction Templates across Different Tasks and Datasets
We provide instruction templates for various graph-related tasks (i.e., node classification, link prediction, and graph-to-text) across datasets. These templates are shown in Table 7.
Table 7. Instruction templates across different tasks and datasets.
Diverse Instructions for Graph Learning
Node Classification (Arxiv)
Input: The compact graph description of this PAPER is listed as follows: Title: {title of the PAPER}. Abstract: {abstract of the PAPER}. Ego graph nodes: {ego graph node list}. One-hop neighbors: {1-hop neighbor list}. Random walks: {random walk paths}.
Standard Instruction: Given the target PAPER with the compact graph description, which of the following subcategories of computer science does this PAPER belong to {the category list}. Directly give the most likely category of this PAPER.
Standard Output: {¡category¿}.
CoT Instruction: Given the classification of target PAPER title with ¡category¿ in the Arxiv dataset, give your explanation based on the provided compact graph description. Focus your analysis on elucidating the reasons behind this classification in a clear Chain of Thought. Keep the analysis brief and to the point.
CoT Output: Considering the PAPER’s Compact graph description, its classification is valid under ¡category¿, because {reasoning process and answers}.
Link Prediction (Arxiv)
Input: The compact graph description of this PAPER 1 is listed as follows: Title: {title of the PAPER 1}. Abstract: {abstract of the PAPER 1}. Ego graph nodes: {ego graph node list}. One-hop neighbors: {1-hop neighbor list}. Random walks: {random walk paths}. The compact graph description of this PAPER 2 is listed as follows: Title: {title of the PAPER 2}. Abstract: {abstract of the PAPER 2}. Ego graph nodes: {ego graph node list}. One-hop neighbors: {1-hop neighbor list}. Random walks: {random walk paths}.
Standard Instruction (pair-wise): Given the compact graph description in PAPER 1 and Paper 2. If the connection between the PAPERs represents the relationship between them, are they connected? Give me a direct answer of ”yes” or ”no”.
Standard Output (pair-wise): {yes} or {no}.
CoT Instruction (pair-wise): Given the established link between PAPER 1 and PAPER 2 in the Arxiv dataset, give your explanation based on the provided compact graph description. Focus your analysis on elucidating the reasons behind this link in a clear Chain of Thought. Keep the analysis brief and to the point.
CoT Output (pair-wise): The connection between PAPER 1 and PAPER 2 seems to be grounded on {reasoning process and answers}.
Graph-to-Text (AGENDA)
Input: The compact graph description of this PAPER is listed as follows: Title: {title of the PAPER}. Ego graph nodes: {ego graph node list}. One-hop neighbors: {1-hop neighbor list}. Random walks: {random walk paths}.
Standard Instruction: Given the target PAPER in the Agenda dataset with the compact graph description. Please generate the target PAPER abstract from the compact graph description.
Standard Output: {abstract of the PAPER} .
CoT Instruction: Given the generated abstract of the target PAPER. Please use the provided compact graph description, to examine how these elements influenced the generation of the abstract with a clear Chain of Thought (CoT). Keep the CoT brief and to the point.
CoT Output: The CoT for this generation is as follows: {the content of the CoT}.

### 13 Mar 2024 v2

@@
 Recently, Graph Neural Networks (GNNs) have become widely used in modeling attributed graphs (hamilton2017inductive, ; wang2019heterogeneous, ; cui2020adaptive, ).
-However, they need to be re-trained whenever applied to different graph datasets and tasks.
+However, they need to be re-trained whenever applied to different graph tasks and datasets.
 Inspired by the great success of Large Language Models (LLMs), the combination of GNNs and LLMs aims to enhance the processing of text-attributed graphs, which can be categorized into two main approaches.
-The first line tries to train GNNs with LLM-enhanced features (e.g., OFA (liu2023one, ), ALL-in-One (sun2023all, ), GIMLET (zhao2023gimlet, ), ExpAsFeat (he2023explanations, ), LLMForGraph (chen2023exploring, ) and PRODIGY (huang2023prodigy, )) The second line tunes LLMs for graph applications (e.g., InstructGLM (ye2023natural, ), NLGraph (wang2024language, ) and LMMoL (qian2023can, ), GPT4GRAPH (guo2023gpt4graph, )).
+The first line tries to train GNNs with LLM-enhanced features (e.g., OFA (liu2023one, ), ALL-in-One (sun2023all, ), GIMLET (zhao2023gimlet, ), ExpAsFeat (he2023explanations, ), LLMForGraph (chen2023exploring, ) and PRODIGY (huang2023prodigy, )).
+The second line tunes LLMs for graph applications (e.g., InstructGLM (ye2023natural, ), NLGraph (wang2024language, ), LMMoL (qian2023can, ), and GPT4GRAPH (guo2023gpt4graph, )).
 Despite the promising direction of integrating LLMs with GNNs, the full exploration of LLMs’ generative capabilities has yet to be harnessed for a wide array of graph-oriented tasks and datasets.
@@
 However, these instructions are often neither readily available nor economically feasible to produce manually.
-While advanced language models like GPT-4 possess the capability to generate diverse instructions, it remains unknown how to effectively leverage these generative abilities to produce relevant, task-specific instructions.
+While advanced language models like GPT-4 possess the capability to generate diverse instructions, it remains unknown how to effectively leverage these generative abilities to produce relevant and task-specific instructions.
 Challenge III: How to properly allocate instructions for graph-oriented instruction tuning?
@@
 However, maintaining broad task and dataset coverage while averting catastrophic forgetting poses a significant challenge, especially when dealing with a variety of tasks and datasets in one generic graph model.
-To tackle these challenges, we propose Graph-oriented Instruction Tuning of Large Language Models for Generic Graph Mining (MuseGraph), which consists of three pivotal steps: (i) Development of Compact Graph Descriptions, where we introduce a novel “node energy” metric to texualize graphs with essential semantic and structural details under limited tokens; (ii) Generation of Diverse Instructions, which invokes the advanced generative abilities of GPT-4 to create CoT-based instruction packages tailored for various graph tasks, thus enriching LLMs’ capabilities in understanding and analyzing graph data without the expense of manual instruction crafting; (iii) Graph-aware Instruction Tuning, which introduces a dynamic instruction package allocation strategy based on the specific needs of each graph mining task, ensuring comprehensive and effective LLM tuning.
+To tackle these challenges, we propose Graph-oriented Instruction Tuning of Large Language Models for Generic Graph Mining (MuseGraph), which consists of three pivotal steps: (i) Development of Compact Graph Descriptions, where we introduce a novel “node energy” metric to textualize graphs with essential semantic and structural details under limited tokens; (ii) Generation of Diverse Instructions, which invokes the advanced generative abilities of GPT-4 to create CoT-based instruction packages tailored for various graph tasks, thus enriching LLMs’ capabilities in understanding and analyzing graph data without the expense of manual instruction crafting; (iii) Graph-aware Instruction Tuning, which introduces a dynamic instruction package allocation strategy based on the specific needs of each graph mining task, ensuring comprehensive and effective LLM tuning.
 Our overall contributions in this work are summarized as follows:
   • Formulation of generic graph mining.
-We establish a generic graph framework that effectively transforms graph structures into LLM-friendly formats while preserving the generative capabilities necessary for diverse graph-oriented tasks
+We establish a generic graph framework that effectively transforms graph structures into LLM-friendly formats while preserving the generative capabilities of necessary for diverse graph-oriented tasks.
   • Effective Model Designs.
-We design and implement a set of models and mechanisms including the development of compact graph descriptions, automatically generating diverse, task-specific instructions, and graph-aware instruction tuning, targeting a generic graph model across tasks and datasets.
+We design and implement a set of models and mechanisms including the development of compact graph descriptions, automatically generating diverse, task-specific instructions, and graph-aware instruction tuning, targeting a generic graph model across different tasks and datasets.
   • Extensive Experiments Across Graph Tasks and Datasets.
@@
 Graph representation learning has emerged as a key technique for the complex structures of networks with abundant attributes (DBLP:journals/debu/HamiltonYL17, ; DBLP:journals/tkde/CaiZC18, ; DBLP:journals/tnn/WuPCLZY21, ; zhou2020graph, ; DBLP:journals/tbd/WangBSFYY23, ).
-Among existing node embedding methods, many had analyzed and utilized the great promise of random walks (RWs) in capturing the topological structures of graphs (perozzi2014deepwalk, ; grover2016node2vec, ; dong2017metapath2vec, ; ivanov2018anonymous, ).
-However, the above methods ignored the abundant attribute information surrounding the nodes and edges (mao2023hinormer, ).
-Recently, Graph neural networks (GNNs) for learning node representations through aggregating information from neighboring nodes on graphs (kipfsemi, ; hamilton2017inductive, ; yang2020relation, ).
-However, most existing GNNs were established in a supervised learning setting, which requires abundant task-specific labeled data that may not be available in real-world applications (zhu2021graph, ; cooray2022graph, ), and the embeddings they learned are not generalizable across different downstream tasks (zhang2018link, ), which had to be re-trained whenever applied to different graph tasks and dataset.
+Among existing node embedding methods, many have analyzed and utilized the great promise of random walks (RWs) in capturing the topological structures of graphs (perozzi2014deepwalk, ; grover2016node2vec, ; dong2017metapath2vec, ; ivanov2018anonymous, ).
+However, the above methods ignore the abundant attribute information surrounding the nodes and edges (mao2023hinormer, ).
+Recently, Graph neural networks (GNNs) learn node representations through aggregating information from neighboring nodes on graphs (kipfsemi, ; hamilton2017inductive, ; yang2020relation, ).
+However, most existing GNNs are established in a supervised learning setting, which requires abundant task-specific labeled data that may not be available in real-world applications (zhu2021graph, ; cooray2022graph, ), and the embeddings that they learned are not generalizable across different downstream tasks (zhang2018link, ), which has to be re-trained whenever applied to different graph tasks and datasets.
 Although some studies tried to reduce the labeling effort by pre-training an expressive GNN model on unlabeled data with self-supervision methods (e.g., contrastive learning) (hu2020gpt, ; zhu2021transfer, ; jiang2021pre, ), their performances in specific downstream tasks still relied much on the properly chosen self-supervision tasks and attribute encoders (zhang2023structure, ; sun2022gppt, ).
@@
 2.2. Leveraging LLMs for Graph Mining
-Figure 2. The overall framework of MuseGraph.
 Table 1.
@@
 OFA (liu2023one, )
-NLGraph (wang2024language, ) $\sim$
-GPT4GRAPH (guo2023gpt4graph, ) $\sim$
-GraphGPT (guo2023gpt4graph, ) $\sim$
-InstructGLM (ye2023natural, ) $\sim$
+NLGraph (wang2024language, )
+$\sim$
+GPT4GRAPH (guo2023gpt4graph, )
+$\sim$
+GraphGPT (guo2023gpt4graph, )
+$\sim$
+InstructGLM (ye2023natural, )
+$\sim$
 MuseGraph
+Figure 2. The overall framework of MuseGraph.
 Motivated by the impressive strides in Large Language Models (LLMs), combining Graph Neural Networks (GNNs) and LLMs is creating substantial progress in handling complex text-attributed graphs (li2024survey, ; chen2023label, ; chen2023exploring, ).
-Methods leveraging this combination fall into two broad categories (shown in Table 1).
+Existing methods leveraging this combination mainly fall into two broad categories (shown in Table 1).
 The first category is training GNNs with LLM-enhanced features, labeled GNN* in Table 1).
@@
 OFA (liu2023one, ) prompted LLMs to translate diverse textual attributes into uniform feature vectors across domains.
-These models primarily utilized LLMs to process textual content within graphs and tend to be tailored to specific domains, often failed to be applied for cross-task and cross-dataset applications and lost the generation capability from LLMs.
+These models primarily utilized LLMs to process textual content within graphs and tend to be tailored to specific domains, while they often failed to be applied for cross-task and cross-dataset applications and lost the generation capability from LLMs.
 Although these models effectively focus on leveraging LLMs’ textual prowess, most of them are domain-specific graph models, which fail to perform cross-task and cross-dataset.
@@
 GraphGPT (tang2023graphgpt, ) with one checkmark on GNN* and a partial checkmark on LLM* jointly trained task-specific GNN and one projector on the top of LLMs, which performed well across different datasets.
-Note that, the partial checkmark for some methods means that they can be extended to cross-dataset targets, but this is not their main research focus and they have no design mechanism to achieve it.
-Compared with GNN*-based methods, LLM*-based methods have the potential to perform generic graph mining with powerful cross-tasks and cross-datasets.
+Note that, the partial checkmark $\sim$  for some methods means that they can be extended to cross-dataset targets, but this is not their main research focus and they have no design mechanism to achieve it.
+Compared with GNN*-based methods, LLM*-based methods have the potential to perform generic graph mining with powerful cross-task and cross-dataset.
 However, such power remains largely under-explored.
@@
 3.1. Overview of Our Framework
-Objective: In this paper, we aim to pursue an ideal graph that can fascinate a more effective and generic approach via graph-oriented instruction tuning of LLMs, through which we wish to seamlessly integrate the strengths of GNNs and LLMs for graph mining across different tasks and dataset
+Objective: In this paper, we aim to pursue an ideal graph that can seamlessly integrate the strengths of GNNs and LLMs via graph-oriented instruction tuning of LLMs, through which we wish to facilitate a more effective and generic approach for graph mining across different tasks and datasets.
 Overview: To achieve this goal, we propose MuseGraph framework, which comprises three major components.
-Firstly, we develop a compact graph description mechanism that captures critical semantic and structural details within the constraints of language tokens.
-Secondly, we generate a diverse range of instructions through the reasoning capabilities of LLMs, thus facilitating task-specific, Chain-of-Thought-based instruction packages for various graph tasks.
+Firstly, we develop a compact graph description mechanism that captures critical semantic and structural details within the constraints of language token limitations.
+Secondly, we generate a diverse range of instructions through the reasoning capabilities of LLMs, thus facilitating task-specific and Chain-of-Thought-based instruction packages for various graph tasks.
 Thirdly, we adopt a graph-aware instruction tuning, utilizing a dynamic allocation strategy for instruction packages tailored to the unique demands of each graph task.
-The overall model architecture is shown in Figure 2, and we elaborate on the main components.
+The overall model architecture is shown in Figure 2, and we elaborate on the three main components.
 3.2. Compact Graph Description
 Leveraging the capabilities of Large Language Models (LLMs) for graphs presents a unique set of challenges, primarily due to LLMs’ inherent limitations in directly inputting graph structures under the token limitation.
-Therefore, it is non-trivial to automatically extract key information from graphs under the token limitations, which require a compact description including complex node and edge attributes along with structural details.
+Therefore, it is non-trivial to automatically extract key information from graphs under the token limitations, which requires a compact description including complex node and edge attributes along with structural details.
 Inspired by common graph analysis techniques such as walks and neighbors, we propose a novel method of textualization to describe graphs via these concepts.
@@
 The integration of neighbors and walks can achieve a holistic understanding of graph structures.
-Given LLM token limits and the varied contribution of neighbors and walks to node understanding, we further develop an adaptive input generation mechanism to ensure the compactness of the description.
-Specifically, we first design a “node energy” metric, $H(v_{i})$ assessing node information from two perspectives: token count in attributes and node degree count.
+Given LLM token limitations and the varied contribution of neighbors and walks to node understanding, we further develop an adaptive input generation mechanism to ensure the compactness of the description.
+Specifically, we first design a “node energy” metric, $H(v_{i})$ assessing node information from two perspectives: token count in node attributes and node degree count.
 This metric enables us to effectively filter and select neighbor nodes and walk nodes, prioritizing those that are abundant in semantic information and possess a significant number of neighboring nodes, thus enhancing the expressiveness of the graph description.
@@
 (1) $H(v_{i})=T(v_{i})*\lceil\log(D(v_{i})+1)\rceil,$
-where $v_{i}$ is the target node, ${T(v_{i})}$ is the number of node tokens processed by a tokenizer and ${D(v_{i})}$ is the number of node degrees.
-10:  Attributed graph $\mathcal{G}$ with $N$ node, token count set $\mathcal{T}$,
+where $v_{i}$ is the target node, ${T(v_{i})}$ is the number of node tokens processed by a tokenizer, and ${D(v_{i})}$ is the number of node degrees.
+10:  Attributed graph $\mathcal{G}$ with $N$ nodes, token count set $\mathcal{T}$,
 node energy set $\mathcal{H}$, target node $v^{*}$, token limitation $L(v^{*})$
@@
 Based on $H(v_{i})$, our method strategically incorporates neighbors and walks, as depicted in Algorithm 1.
-A neighbor, $v_{i}$, is chosen to describe the target $*v*$ if its $H(v_{i})$ surpasses $H(v^{*})$, ensuring the included node can provide supplemental information.
+A neighbor $v_{i}$ is chosen to describe the target $v^{*}$ if its $H(v_{i})$ surpasses $H(v^{*})$, ensuring the included node can provide supplemental information.
 The process of including walks concludes when encountering a node whose $H(v_{i})$ does not meet the threshold set by the target’s $H(v^{*})$, thus refining the input to maximize relevance within the constraints of the token limit.
-To further demonstrate the adaptive input generation for neighbors and walks tailored to each node, we present an example in Figure 3).
+To further demonstrate the adaptive input generation for neighbors and walks tailored to each node, we present an example in Figure 3.
 Node $v_{1}$’s with a relatively low $H(v_{1})$ value, includes a wide range of neighbors to capture more context, excluding $v_{4}$ due to its even lower $H(v_{4})$.
@@
 This frees up tokens to detail more walks for $v_{2}$.
-The $H$ metric thus finely balances neighbor and walk inclusion for each node, marrying information-rich and token-efficient.
+The $H$ metric thus finely balances neighbor and walk inclusion for each node, marrying information-rich and token-efficient characteristics.
 The process culminates in the textualization of each node’s key information, producing a tailored and compact graph description as depicted in the upper right of Figure 2.
@@
 Figure 4. A process showing how to distill capabilities from GPT-4 via Chain-of-Thought with node classification.
-Specifically, we first design diverse CoT templates for targeted popular graph tasks, such as node classification, link prediction, and graph-to-text (shown in A Table 7).
+Specifically, we first design diverse CoT templates for targeted popular graph tasks, such as node classification, link prediction, and graph-to-text (shown in Appendix A.3 Table 7).
 Then, we prompt GPT-4 to generate a small number of CoT-based instructions across diverse tasks.
@@
 LLaMA) (touvron2023llama, ) with LoRA (hu2021lora, ) as our starting point for fine-tuning.
-Then, based on diverse instruction set =$\mathcal{I}=\{I_{1},I_{2},\dots,I_{N}\}$ as model input, we adopt the negative log-likelihood loss as the fine-tuning objective (achiam2023gpt, ; ye2023natural, ), which is formulated as follows:
+Then, based on diverse instruction set =$\mathcal{I}=\{I_{1},I_{2},\dots,I_{D}\}$ as model input, we adopt the negative log-likelihood loss as the fine-tuning objective (achiam2023gpt, ; ye2023natural, ), which is formulated as follows:
 (2) $p_{\theta}\left(Y_{j,k}|I_{j},Y_{j,<k}\right)=LLM_{\theta}\left(I_{j},Y_{j,<k}\right),$
@@
 4.1. Experimental Setup
-4.1.1. Datasets Descriptions.
+4.1.1. Datasets.
 To comprehensively verify the effectiveness of our method, we use three real-world datasets for node classification, i.e., OGB-arxiv (abbr.
@@
 4.2. Overall Performance Comparison (RQ1)
-To evaluate our MuseGraph framework, we conducted extensive testing across node classification and graph-to-text tasks using five real-world datasets.
+To evaluate our MuseGraph framework, we conduct extensive testing across node classification and graph-to-text tasks using five real-world datasets.
 4.2.1. Main Results Compared with GNN-based Methods.
@@
 In particular, GLNN is optimized for attribute-rich datasets (e.g., Arxiv and Cora) through its knowledge distillation mechanisms, while GAT is tailored for datasets with basic attributes (e.g., MIMIC-III) through its ability to weigh nodes differently.
-Compared with GLNN and GAT, our MuseGraph with graph-ware instruction tuning with the diverse CoT-based instruction packages can effectively comprehend graph data and achieve accurate predictions of various downstream tasks.
+Compared with GLNN and GAT, our MuseGraph with graph-aware instruction tuning with the diverse CoT-based instruction packages can effectively comprehend graph data and achieve accurate predictions of various downstream tasks.
 One step further, since one key challenge of node classification lies in the generalizability and adaptability of models, we design a few-shot setting on Cora to evaluate models in extending knowledge to unseen scenarios and adapting to new tasks with limited training data.
@@
 MuseGraph, with less pre-training data towards graph-to-text task and mainly targeting generic graph mining, still retains competitive performance against these models by effectively leveraging our graph-oriented instruction tuning of LLMs.
-Note that, MuseGraph can sometimes exceed all baselines with ROUGE-L metric, achieving an average performance gain of 22.40% over the second-runner T5 on average.
-By extracting the compact graph description by adaptively allocation the neighbors and walks under the input token limitation, MuseGraph adeptly extracts a compact graph description.
+Note that, MuseGraph can sometimes exceed all baselines with ROUGE-L metric, achieving an average performance gain of 1.58% over the second-runner T5 on average.
+By adaptively allocating the neighbors and walks under the input token limitation, MuseGraph adeptly extracts a compact graph description.
 This enables it to accurately grasp both the semantic and structural nuances of graphs through a graph-aware fine-tuning mechanism.
@@
 To study the effectiveness of different allocations of CoT-based instruction packages across various tasks and datasets, we conduct ablation studies for node classification on Arxiv, which is based on our proposed compact graph description for inputting graph.
-Using LLaMA as our baseline model, the studies evaluated our MuseGraph across various configurations, while maintaining an equal total count of instruction packages.
+Using LLaMA as our baseline model, the studies evaluate our MuseGraph across various configurations, while maintaining an equal total count of instruction packages.
 We study our MuseGraph as follows:
-  • 1 Task+1 Dateset: Utilizes our proposed CoT-based instructions for node classification on Arxiv.
-  • 2 Task+1 Dateset: Incorporates instructions from both node classification and link prediction on Arxiv.
-  • 2 Task+2 Dateset: Expands to include node classification and link prediction instructions from both Arxiv and Cora.
-  • 3 Task+3 Dataset: Engages with instructions across node classification, link prediction on Arxiv and Cora, and graph-to-text on AGENDA..
+  • 1 Task+1 Dataset: Utilizes our proposed CoT-based instructions for node classification on Arxiv.
+  • 2 Task+1 Dataset: Incorporates instructions from both node classification and link prediction on Arxiv.
+  • 2 Task+2 Dataset: Expands to include node classification and link prediction instructions from both Arxiv and Cora.
+  • 3 Task+3 Dataset: Engages with instructions across node classification, link prediction on Arxiv and Cora, and graph-to-text on AGENDA.
 As shown in Figure 5, we have the following observations:
@@
 This indicates that reasonable allocation of the different volumes of CoT-based instruction packages across datasets can help LLMs effectively and efficiently perform predictions of graph-related tasks.
-With adding more types of tasks and datasets, our proposed MuseGraph can achieve further improvements, where 3 Task+3 Dataset overperforms 2 Task+2 Dataset ranges from 3.03% in Macro-F1 to 3.94% in Weighted-F1.
+With adding more types of tasks and datasets, our proposed MuseGraph can achieve further improvements, where 3 Task+3 Dataset outperforms 2 Task+2 Dataset ranging from 3.03% in Macro-F1 to 3.94% in Weighted-F1.
 Such observation strongly indicates that the dynamic instruction package allocation strategy can enhance the accuracy of graph tasks and avert catastrophic forgetting, where LLMs might regress in their generative abilities.
 4.4. Case Studies (RQ3)
-To demonstrate the advantages of MuseGraph over the LLM-based methods on the generation tasks, we provide the abstract of graph-to-text task prompted by LLaMa, GPT-3.5, and our MuseGraph on AGENDA.
+To demonstrate the advantages of MuseGraph over the LLM-based methods on the generation tasks, we provide the abstract of the graph-to-text task prompted by LLaMA, GPT-3.5, and our MuseGraph on AGENDA.
 The provided graph structure and its ground-truth abstract are displayed at the top of Table 6.
-We standardized the process by employing our compact graph description as the input across all models.
+We standardize the process by employing our compact graph description as the input across all models.
 As shown in Table 6, the abstract generated by LLaMA offers a basic summary while overlooking some key details, such as the specific incremental learning aspect of the fuzzy algorithm (shown in the blue text).
-Compared with LLaMA, Chatgpt-3.5 generates even three times longer content with an elaborate description.
+Compared with LLaMA, GPT-3.5 generates even three times longer content with an elaborate description.
 However, the model inevitably includes some fabricated sentences, such as the “Through extensive experiment and evaluation, …” in red, which is irrelevant to the provided ground-truth abstract.
 Since our method performs graph-aware instruction tuning with compact graph descriptions, we can produce precise abstracts by extracting the key information from the graph.
-Moreover, by incorporating the distilled reasoning and generative capacity from CoT-based instructions, MuseGraph can not only produce an abstract that aligns with the core details of the ground-truth but also provides fluent and precise CoT inherited the advanced large language models.
+Moreover, by incorporating the distilled reasoning and generative capacity from CoT-based instructions, MuseGraph can not only produce an abstract that aligns with the core details of the ground truth but also provides fluent and precise CoT inherited from the advanced large language models.
 5. Conclusion
@@
   • LLaMA-V1-7B [53] is an open-sourced Large Language Model (LLM) designed for natural language understanding and generation tasks.
-  • ChatGPT-3.5 [41] uses Graph Neural Networks (GNNs) and transformer-based models to generate coherent text based on input structured graphs.
+  • GPT-3.5 [41] uses both Graph Neural Networks (GNNs) and transformer-based models to generate coherent text based on input structured graphs.
 A.3. Instruction Templates across Different Tasks and Datasets
@@
 Random walks: {random walk paths}.
-Standard Instruction: Given the target PAPER with the compact graph description, which of the following subcategories of computer science does this PAPER belong to {the category list}.
+Standard Instruction: Given the target PAPER with the compact graph description in the Arxiv dataset, which of the following subcategories of computer science does this PAPER belong to {the category list}.
 Directly give the most likely category of this PAPER.
@@
 Abstract: {abstract of the PAPER 1}.
-Ego graph nodes: {ego graph node list}.
-One-hop neighbors: {1-hop neighbor list}.
-Random walks: {random walk paths}.
+Ego graph nodes: {ego graph node list of the PAPER 1}.
+One-hop neighbors: {1-hop neighbor list of the PAPER 1}.
+Random walks: {random walk paths of the PAPER 1}.
 The compact graph description of this PAPER 2 is listed as follows: Title: {title of the PAPER 2}.
 Abstract: {abstract of the PAPER 2}.
-Ego graph nodes: {ego graph node list}.
-One-hop neighbors: {1-hop neighbor list}.
-Random walks: {random walk paths}.
-Standard Instruction (pair-wise): Given the compact graph description in PAPER 1 and Paper 2.
+Ego graph nodes: {ego graph node list of the PAPER 2}.
+One-hop neighbors: {1-hop neighbor list of the PAPER 2}.
+Random walks: {random walk paths of the PAPER 2}.
+Standard Instruction (pair-wise): Given the compact graph descriptions of PAPER 1 and PAPER 2 in the Arxiv dataset.
 If the connection between the PAPERs represents the relationship between them, are they connected?
-Give me a direct answer of ”yes” or ”no”.
+Give me a direct answer of “yes” or “no”.
 Standard Output (pair-wise): {yes} or {no}.
@@
 Standard Output: {abstract of the PAPER} .
-CoT Instruction: Given the generated abstract of the target PAPER.
+CoT Instruction: Given the generated abstract of the target PAPER in the AGENDA dataset.
 Please use the provided compact graph description, to examine how these elements influenced the generation of the abstract with a clear Chain of Thought (CoT).